# Effective Multi-Dimensional 3D Scatterplots as 2D Figures

### v2

@@ IV-B Software infrastructure
Alongside the script, the server hosts the 198 images in a subfolder. The script injects those images into the HTML code using the base64 format. This is done in order to minimise the possibility of giving more information based on the image filename, or enabling the crawling of all the images. These images have been generated in a semi-automated way, as the current software is GUI-focused. Another Python script has been used to generate the 6 random datasets, and then the GUI was used (incrementally) to export the images. The marking of the random points is done manually using GIMP relatively quickly by opening all images in tabs at once. Finally, the images are overwritten in the .png format (Portable Network Graphics), though a separate bash script compresses them into the more modern and space-efficient .webp format (WebP). The resulting images have a resolution of 1700 by 1275, and consume up to 60 KB each, which also eases their serving through HTML injection. A non data-collecting version of the quiz is kept online [^2] Available after publication..
IV-C Participation
@@ IV-D Results
Figure 8: Mean distance and median time taken for each question type.
@@
Normalisation: In order to remove potential biases, such as with a variation of the user skill or the difficulty of the random dataset, we apply normalisation at different levels. The goal is to increase the reliability of the time/distance comparison especially for more subtle observations.
@@
Figure 9: Study results after applying normalisation.
@@
Most of the observations made on the non-normalised results (figure 8) still hold with normalisation. F (projections) stops being competitive, and there are more visible trade-offs in the center points (outside the Pareto front of figure 9). This shows the impact of normalisations, but also overall verifies the well-roundness of the results, as with the sample size and questions. A more subtle finding is about the pairs that only differ in time. In the first instance, F and G (non-coloured and depth-coloured projections respectively) perform similarly in terms of accuracy, but the colouring could have caused some delay from being an unfamiliar approach. In the other instance, C and E are the respective figures without the projections, but here the colour has reduced the time required to do the reading. Overall, there are clear winners, but for points with trade-offs, the unfamiliarity or information crowding could be a notable factor for time required to make the readings.
Pairwise comparison: The colour-mapped matrices of figure 10 can be used to make direct comparisons from question type to question type. The question types are also sorted according to the corresponding performance metric (distance for the top, time for the bottom), also denoting a ranking. The two rankings are relatively similar permutations of the types, with G (colour-mapped projections) being displaced the most, from 6th for distance to 9th for time.
Figure 10: Comparing the performance of question types in pairs.
If we select the best performing (K, three projections/light paths), it can be seen that it provides 5.18x lower distance and 1.66x lower time than the frequently-used isometric plot without additional cues (A). Additionally, incremental changes can be compared directly including the FoV variations, as with 75 (C) over isometric (A) that reduces error by 1.41x and time by 1.09x. It is also interesting to observe the colour clustering, such as with K, J, H and I which perform the best in terms of distance, as they include light paths that point to at least one pane (2-axis combination) for guidance.
@@
The main observation from the study is that the most effective technique is the light paths with projections (types J and K, with 2 and 3 light paths respectively). Another important finding relates to the FoV, as the popularly-used isometric axes give unacceptable performance, while the higher FoV value (75) gives the most promising results among the types comparing FoV values. Other less-conventional features such as fog, depth-based point-colouring and surface ticks have not caused a substantial improvement as additional visual cues, but this could also partly be attributed to the lack of training.
V Related work
The literature can be divided into visualisation frameworks and user studies. The first are a selection of software packages having a 3D scatterplot functionality, including with special visual cues, and can be seen as competing or alternative approaches to the features of the proposed framework. The user studies are on the topics discussed in the paper such as depth perception and plot comprehension, and include empirical results.
V-A Frameworks
With respect to the dimension assignment facility with checkboxes (as in figure 4), it exists in certain software with a limited dimensionality. One instance is “DB Browser for SQLite”, an open source database management system focusing on local databases [15]. It also operates on tabular data, though its functionality is restricted to 2D plots. ParaView [16] is an advanced visualisation framework that also provides point-and-click functionality for the creation of plots. For more advanced figures such as 3D scatterplots with point projections, additional coding is required with a scripting language.
Other examples of software for generating advanced 3D scatterplots include pygmt.Figure.plot3d [17] and the research-informed ParaProf [18] that also uses Java. Both are relatively simplistic when compared to the presented framework, and lack the discussed depth perception and label legibility functionality. Another scatterplot solution that uses JavaFX 3D and the FXyz3D component library is within the deep space trajectory explorer [19], which specialises on multibody trajectory analysis and is not publicly available.
Basic point projections on the back of the view box (similar to the floor of figure 4 (right)) is an existing approach for improving 3D plot readability. One example is a WebGL online 3D scatterplot viewer [20] that features point projections, as well as drop lines when the user hovers the mouse cursor over a point. When compared to the presented software, it is mostly an interactive proof-of-concept, with same-sized points and labels lacking an automatic heading adjustment.
A close framework in regard to the explored functionality is OriginPro [21]. It supports a superset of plot types, but its 3D scatterplots lack the aforementioned unique features. Underneath, a mix of 2D raster (renders) and vector components are used. This creates an impressive clarity effect in printouts (axis labels in a text format), but also results in some loss of physical properties and navigation flexibility (such as quantised rotation ranges). The visual quality is, hence, variable, as with pixellation in the grid lines when zooming in. There is a FoV adjustment support, though limited to 10 increments. Being a more general and proprietary analytics framework, this is reflected in the license cost (\$645 for a node-locked annual subscription).
The presented features are to the best of our knowledge unique (as introduced in the list of section I), though keeping intuitiveness in mind, it could still be possible to find resemblances in related work, or be able to reproduce them with additional scripting. Thus, the contribution in this aspect is at least their coexistence in the proposed open-source framework.
V-B Empirical studies
VI Discussion
This section discusses potential future work including with regard to the functionality of the software, as well as limitations of the research overall including the empirical study.
VI-A Future work
A feature that would be appropriate for future consideration is the increase of the number of available dimensions, such as when involving timestamps [26]. A video or animation generation approach could be developed that linearises one of the dimensions to facilitate the smooth evolution of the dataset across time. Additionally, the amount of realism could be increased to further improve the depth perception and visual cues for more accurate and faster readings. This can include ray-tracing, stereoscopic view for virtual reality (not currently supported in JavaFX), and more complex illumination [27].
In order to increase the generality and adaptability of the framework, conventional features of plotting and data analysis software could be worthwhile to adopt. These include a scripting language for automation, increasing the number of data sources, as with live feeds of streaming data, and error bars for our 3D points. Additional effort can be dedicated to directly distributing the software through open-source software repositories and app stores.
An analogous empirical user study could elaborate on the intuitiveness and effectiveness of 3D scatterplots using additional functionality, as well as the interface itself.
VI-B Limitations
A technical limitation in the framework is with JavaFX that does not support transparency for overlapping objects. Specifically, transparency is one-dimensional in JavaFX meaning that the object that is added to the world later enables seeing past objects through it, but not the other way around. As a workaround, when adjusting the viewing angle inside the GUI, our framework sorts the visible objects according to the distance from the camera. This achieves the required result, which is to be able to see the objects behind a transparent object correctly. However, when the scene is complex enough, such as an object spanning across multiple other objects, the plot may lose some of its realism.
@@ VI-B Limitations
VII Conclusions
In this paper, a novel open-source framework is presented for plotting multi-dimensional 3D scatterplots. It provides an intuitive design as a JavaFX desktop application for quickly applying advanced visual cues and additional dimensions to 3D scatterplots. A series of features relate to the easy association and representation of multiple dimensions from tabular data, as well as to the increase of depth perception. The framework is indicatively demonstrated with the use of recent datasets on high-performance computing and the world happiness report, where a series of insights are drawn from single multi-dimensional plots.
The empirical study from the 57 participants shows that certain visual cue combinations can dramatically affect the effectiveness of 3D scatterplots. When comparing to a conventional plot with isometric axes, the best candidate (wider field of view, 3 light paths and projections) reduces the read error (distance) for discrete data by 5.2 times on average, while also reducing the reading time by 1.7 times. Simpler adjustments also resulted in conclusive performance results, such as with the exaggerated field of view of 75 which reduces the mean distance and time taken by 1.4 and 1.1 times respectively.

### v3

@@ 3.2 Software infrastructure
Alongside the script, the server hosts the 198 images in a subfolder. The script injects those images into the HTML code using the base64 format. This is done in order to minimise the possibility of giving more information based on the image filename, or enabling the crawling of all the images. These images have been generated in a semi-automated way, as the current software is GUI-focused. Another Python script has been used to generate the 6 random datasets, and then the GUI was used (incrementally) to export the images. The marking of the random points is done manually using GIMP relatively quickly by opening all images in tabs at once. Finally, the images are overwritten in the .png format (Portable Network Graphics), though a separate bash script compresses them into the more modern and space-efficient .webp format (WebP). The resulting images have a resolution of 1700 by 1275, and consume up to 60 KB each, which also eases their serving through HTML injection. A non data-collecting version of the quiz is kept online.
The software that is used to generate the 3D scatterplots is developed in-house, and is open source [14]. It is written in Java and uses JavaFX and FXyz for the 3D functionality. The plots and their labels exist as 3D models in their entirety, which enables the support of all of the aforementioned visual cues.
3.3 Participation
@@ 4 Results
Figure 4: Mean distance and median time taken for each question type
4.1 Basic insights
@@
4.2 Normalisation
In order to remove potential biases, such as with a variation of the user skill or the difficulty of the random dataset, we apply normalisation at different levels. The goal is to increase the reliability of the time/distance comparison especially for more subtle observations.
@@ 4.2 Normalisation
Figure 5: Survey results after applying normalisation
@@
After considering the last observations, most still hold with normalisation. F (projections) stops being competitive, and there are more visible trade-offs in the center points (outside the Pareto front). This shows the impact of normalisations, but also overall verifies the well-roundness of the results especially on the sample size and questions. A more subtle finding is about the pairs that only differ in time. In the first instance, F and G (non-coloured and depth-coloured projections respectively) perform similarly in terms of accuracy, but the colouring could have caused some delay from being an unfamiliar approach. In the other instance, C and E are the respective figures without the projections, but here the colour has reduced the time required to do the reading. Overall, there are clear winners, but for points with trade-offs, the unfamiliarity or information crowding could be a notable factor for time required to make the readings.
5 Discussion
This section overviews the key findings from the survey with the help of pairwise comparisons, while elaborating on potential limitations and future work on the methodology.
5.1 Key survey insights
@@
The main observation from the survey is that the most efficient technique is the light paths with projections (types J and K, with 2 and 3 light paths respectively). These correspond to the hypotheses H4 on projections, H6 on light paths, and H7 on combinations, as introduced in section 1, and are all confirmed. Another important finding relates to the FoV, as the popularly-used isometric axes gave unacceptable performance, while the higher FoV value (75) gave the most promising results among the types comparing FoV values. Therefore, H1 on exaggerated FoV is mostly supported, even though the reduced variation when having a moderate FoV is notable. Other less-conventional features such as fog, depth-based point-colouring and surface ticks have not caused a substantial improvement as additional visual cues, but this could also partly be attributed to the lack of training. The latter three correspond to hypotheses H2 on fog, H3 on colour-mapping and H5 on surface ticks, which are not supported in this exploration.
Figure 6: Comparing the performance of question types in pairs
The colour-mapped matrices of figure 6 can be used to make direct comparisons from question type to question type. The question types are also sorted according to the corresponding performance metric (distance for the top, time for the bottom), also denoting a ranking. The two rankings are relatively similar permutations of the types, with G (colour-mapped projections) making the most steps in the ranking, from 6th to 9th for time. If we select the best performing (K, three projections/light paths), it can be seen that it provides 5.18x lower distance and 1.66x lower time than the frequently-used isometric plot without additional cues (A). Additionally, incremental changes can be compared directly including the FoV variations, as with 75 (C) over isometric (A) that reduces error by 1.41x and time by 1.09x. It is also interesting to observe the colour clustering, such as with K, J, H and I which perform the best in terms of distance, as they include light paths that point to at least one pane (2-axis combination) for guidance.
5.2 Limitations
@@
6 Conclusions
This study shows that certain visual cues can dramatically affect the effectiveness of 3D scatterplots. An online survey requested 57 participants from an academic community to provide readings from marked 3D points. When comparing to a conventional plot with isometric axes, the best candidate (wider field of view, 3 light paths and projections) reduces the read error (distance) for discrete data by 5.2 times on average, while also reducing the reading time by 1.7 times. Simpler adjustments also resulted in conclusive performance results, such as with the exaggerated field of view of 75 which reduces the mean distance and time taken by 1.4 and 1.1 times respectively. A series of measures has been applied before and after the data collection to ensure the impartiality and appropriateness of the results. These include the uniqueness between survey questions and instances, and the normalisation steps on the user competency and the difficulty per random dataset. IO